OpenStudy (anonymous):

How many pairs of these numbers add up to a multiple of 11. 9, 46, 79, 13 64, 90, 2, 97, 25, 31, 20, 22 4, 52, 55, 7 How could you convince someone that they haven't missed any? Notice anything interesting about the pairs of numbers? Can you explain what's happening?

OpenStudy (hba):

Yeah @chrisabraham87

OpenStudy (amistre64):

a + b = 0 (11) a = -b (11) maybe

OpenStudy (anonymous):

Any ideas?

OpenStudy (hba):

A grid.

OpenStudy (anonymous):

Can you answer the questions please?!

OpenStudy (hba):

Sure just let me figure it out.

OpenStudy (anonymous):

Thank you

OpenStudy (hba):

Follow the numbers and see where they go.

OpenStudy (anonymous):

?

OpenStudy (amistre64):

chris, please dont expect us to answer your question for you without you providing some input and effort towards a solution.

OpenStudy (anonymous):

That's 1 pair. ------------ Remove 22 & 55 ----- 9, 46, 79, 13 64, 90, 2, 97, 25, 31, 20, 4, 52, 7 14 numbers remain. Try them 2 at a time, starting with 9. There are 5 pairs with 9. Eliminate the 9, try 46. 46, 79, 13 64, 90, 2, 97, 25, 31, 20, 4, 52, 7 There are 4 more pairs with 46. Eliminate the 46, etc. 79, 13 64, 90, 2, 97, 25, 31, 20, 4, 52, 7 4 pairs with 79 ----------- 13 64, 90, 2, 97, 25, 31, 20, 4, 52, 7 4 pairs with 13 -------------- 64, 90, 2, 97, 25, 31, 20, 4, 52, 7 2 pairs with 64 ------------- 90, 2, 97, 25, 31, 20, 4, 52, 7 3 with 90 ---------- 2, 97, 25, 31, 20, 4, 52, 7 3 with 2 ----------- 97, 25, 31, 20, 4, 52, 7 None with 97 ------------ 25, 31, 20, 4, 52, 7 1 w/25 ------------ 31, 20, 4, 52, 7 None w/31 ------ 20, 4, 52, 7 None w/20 ----------- 1 w/4 and that's all 28 total.

OpenStudy (anonymous):

There are 16 numbers. 2 of them are multiples of 11, 22 & 55, so they can only be added to each other --> 77. Adding any other number to them won't work.

OpenStudy (amistre64):

oy thats alot to read :) good job tho

OpenStudy (anonymous):

Posts should be other way around. Can anyone help me answering the 3 questions on the task?

OpenStudy (amistre64):

is the block of numbers a single set? or is that 4 different sets?

OpenStudy (anonymous):

It's a grid

OpenStudy (anonymous):

So any combination of numbers so long as they are multiples of 11

OpenStudy (amistre64):

is this number theory by chance?

OpenStudy (anonymous):

I am not sure at all

OpenStudy (amistre64):

no idea on the first question, i cant even convince some people that 1+1 =2 ....

OpenStudy (amistre64):

the smallest pair, im assuming a pair to be the sum of 2 number, is 2+4=6 lhe largest pair is 97+90 = 187 to give a range of values between 6 and 187

OpenStudy (anonymous):

Must be multiples of 11 mate

OpenStudy (amistre64):

i know that multples of 11 in this range take the form: 11(a,b) = a, a+b, b 11(n) = 187 ; therefore n=17 right, nust be a multiple of 11, and since i dont know a cookie cutter response there is a systematic process to go thru to determine if a suitable solution exists

OpenStudy (amistre64):

ive never seen the grid tho, so i cant say i know if this pertains to number theory or something else

OpenStudy (amistre64):

11 = 10+1 11n = 10n + n 79+25 = (70+20) + (9+5) = (70+20+10) + (4) = 10(10) + 4 ; since 10 not equal 4, this isnt a multiple

OpenStudy (anonymous):

alright, here's the short way to do this: rewrite all the numbers in mod 11 (i.e. remainder when divided by 11). you get: 9,2,2,2,9,2,2,9,3,9,9,0,4,8,7 that's five 2's, five 9's, one 3, one 8, one 4, one 7. any of the 2's paired with any of the 9's gets u 11, counting that gives 25 pairs, then the 3-8 is one pair, the 4-7 is one pair, for a grand total of 27 pairs

OpenStudy (anonymous):

oops, forgot 22 and 55, so that's one more pair. grand total: 28 pairs

OpenStudy (amistre64):

believe it or not, i just logged on to suggest the same thing :)